OpenStudy (anonymous):

f(x) = 5/(7x-8) What is the inverse of this function?

OpenStudy (asnaseer):

you need to re-arrange this equation so that you get x in terms of f(x). that will then be your inverse function. e.g. suppose we have f(x) = x + 1 then we can re-arrange this to get x = f(x) - 1 so the inverse function is:\[f^-1(x)=x-1\]

OpenStudy (asnaseer):

the last line should be:\[f^{-1}(x) = x-1\]

OpenStudy (anonymous):

\[y=\frac{5}{7x-8}\]Exchange the x and y variables\[x=\frac{5}{7y-8}\]solve for y (cross multiply)\[x(7y-8)=5\]\[7y-8=\frac{5}{x}\]Add 8\[7y=\frac{5}{x}+8\]Get a common denominator\[7y=\frac{5+8x}{x}\]multiply by 1/7\[y=\frac{5+8x}{7x}\]this is the inverse function

OpenStudy (anonymous):

One characteristic of a function and its inverse is that they are symmetric about the line y=x